OpenStudy (anonymous):

12 is what percent of 75? 0.16% 6.25% 9% 16%

OpenStudy (anonymous):

0.16%

OpenStudy (anonymous):

how do you get it?

OpenStudy (anonymous):

xis the unknown percent. \[ \frac{x}{100} * 75 = 12 \] solve for x.

OpenStudy (anonymous):

sorry i think its 16% not 0.16%

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Solve for x and you'll know what it is.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

multiply both sides by 100 / 75 and see what happens.

OpenStudy (anonymous):

\[ \frac{x}{100}*75 * \frac{100}{75} = 12 * \frac{100}{75} \]